I am a parent trying to help my 8th grader. Could you please help me explain how to work this equation: 4(5c+3)-3(4c-6)=110......I think it would be
20c+12-12c+18=110 but then what do we do?

Answer by Linz(9)

If you are trying to find the value of c:
You did the first part right. Now you need to combine like terms.
20c-12c=8c
18+12=30
(your equation should look like this)
8c+30=110
then we need to isolate the variable,so..
subtract 30 from both sides of the equal sign.
8C=80
then divide both sides by 8
80 divided by 8 is 10
So c=10
Hope this helps!!
Linz